OpenStudy (anonymous):

Can someone show me the steps? Factor completely: 3x2 + 12x + 7

OpenStudy (hba):

did u try to solve it

OpenStudy (anonymous):

use quadratic equation where a=3 b=12 and c=7 place them equation nad you have your answer =)

OpenStudy (anonymous):

Can you show me the steps?

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

is this the answer: (3x+7)(x+1) ?

OpenStudy (anonymous):

\[-(12)\pm \sqrt{12^2-4(3)(7)}/2(3)\] \[-(12)\pm \sqrt{144-84}/(6)\] \[-(12)\pm \sqrt{60}/(6)\] \[-(12)\pm 7.75/6\] \[-(12)- 7.75/6\]=-3.29 \[-(12)+ 7.75/6\]=-0.7

OpenStudy (anonymous):

sorry for replying you so late actually my internet is not working properly :(